# Data-adaptive structural change-point detection via isolation

### v1

@@ 2.2 The DAIS algorithm
In the following paragraphs, the method is presented in more generality. For the rest of the paper, we use $s,e$ to denote the start- and the end-point, respectively, of the interval under consideration, while $T\in\mathbb{N}$ denotes the total length of the given signal. $N\in\mathbb{N}_{0}$ and $r_{i}$ for $i=1,2,\dots,N$ are the number and locations of the true change-points sorted in increasing order, while $\hat{N}$, $\hat{r}_{i}$ for $i=1,2,\dots,\hat{N}$ are their estimated values. We denote by $\lambda_{T}\in\mathbb{N}$ the expansion parameter and $\zeta_{T}\in\mathbb{R^{+}}$ the threshold, which is used to decide if a point is a change-point; more details on their choices are given in Section 4.2. In addition, $d_{s,e}$ is the location of the largest difference in the interval $[s,e]$ and is defined as
@@
We denote by $C^{b}_{s,e}(\boldsymbol{X})$ the contrast function applied to the location $b$ using the values $\{X_{s},X_{s+1},\ldots,X_{e}\}$ for $b\in[s,e)$. The functions used for the cases of piecewise-constant and continuous, piecewise-linear signals are explained in Sections 3.1 and 3.2, respectively. For $K^{l}=\lceil\frac{d_{s,e}-s+1}{\lambda_{T}}\rceil$, $K^{r}=\lceil\frac{e-d_{s,e}+1}{\lambda_{T}}\rceil$, $K^{\max}=\max\{K^{l},K^{r}\}$, $K^{\min}=\min\{K^{l},K^{r}\}$ and $K=K^{\max}+K^{\min}$ the start- and end-points of the intervals $(c_{m}^{l},c_{k}^{r})$ in which the algorithm tests for the existence of a change-point, are given by
@@
These are collected in the ordered sets, $L^{b}_{d_{s,e},s,e}$ and $R^{b}_{d_{s,e},s,e}$, defined in (2.2), which consist of the left and right end-points, respectively, of the intervals used in DAIS. Since the expansions occur in turn once from the right and once from the left, the points appear twice in each set, up until the moment when one end-point is equal to either $s$ or $e$. The left- and right-expanding intervals checked are collected in the ordered set $I_{j}$ in (4). The ordered sets of end-points for the expansions are:
$L^{b}_{d_{s,e},s,e}=\bigl{\{}c^{l}_{0},c^{l}_{1},c^{l}_{1},c^{l}_{2},\dots,c^{l}_{K^{\min}-1},c^{l}_{K^{\min}-1},c^{l}_{K^{\min}},c^{l}_{K^{\min}+1},\ldots,c^{l}_{K^{\max}}\bigr{\}},$
$R^{b}_{d_{s,e},s,e}=\bigl{\{}c^{r}_{1},c^{r}_{1},c^{r}_{2},c^{r}_{2},\dots,c^{r}_{K^{\min}},c^{r}_{K^{\min}},c^{r}_{K^{\min}+1},c^{r}_{K^{\min}+2},\ldots,c^{r}_{K^{\max}}\bigr{\}},$ (3)
@@
$I_{j}=\bigl{[}s_{j},e_{j}\bigr{]}=\Bigl{[}L^{b}_{d_{s,e},s,e}[j],R^{b}_{d_{s,e},s,e}[j]\Bigr{]},\textrm{ for }j\in\{1,2,\ldots,K\},$ (4)
where A[i] denotes the $i^{th}$ element of the ordered set A. It is easy to show that $\lvert L^{b}_{d_{s,e},s,e}\rvert=\lvert R^{b}_{d_{s,e},s,e}\rvert=K$, where $\lvert A\rvert$ denotes the cardinality of the set A. Note that the largest number of expansions required in order to check the whole length of the interval $[s,e]$ around $d_{s,e}$ is $K=\lceil\frac{e-s+1}{\lambda_{T}}\rceil+1$.
@@
     $b_{i}=\rm{argmax}_{t\in[s_{i},e_{i})}C_{s_{i},e_{i}}^{t}\left(\textbf{X}\right)$
     if $C_{s_{i},e_{i}}^{b_{i}}\left(\textbf{X}\right)>\zeta_{T}$ then
@@
Algorithm 1 provides a pseudocode that briefly explains DAIS’ workflow. The first step of DAIS is to identify, within the observed data, $X_{t}$, $t=1,2,\ldots,T$, the location of the largest difference $d_{1,T}$, as defined in (2). In the case that $f_{t}$ is piecewise-constant, $d_{1,T}$ is the value of $t$ that maximizes the pairwise absolute differences between consecutive observations. In the case that $f_{t}$ is continuous, piecewise-linear, $d_{1,T}$ is the value of $t$ that maximizes the absolute value of the twice differenced data. Applying the appropriate transformation on the unobserved true signal $f_{t}$ in (1), will reveal a constant signal at 0, with non-zero values at the locations where the change-points occur, so we expect that the largest difference of $X_{t}$ will be at a point near the change-point with the jump of the largest magnitude. A proper explanation of what we consider to be ‘near’ is discussed in Section 2.3 in detail. After calculating $d_{1,T}$, expanding intervals are used to check for possible change-points. The algorithm expands the intervals once from the right and once from the left at every step by an expansion parameter of magnitude $\lambda_{T}$. This is done in order to identify the location of the change-point that is closest to the detected largest difference, which is the reason why the algorithm is data-adaptive. This expansion is performed for at most $\lceil\frac{T}{\lambda_{T}}\rceil+1$ steps in total. Upon the detection of a change-point, the algorithm restarts after discarding the data that have already been checked. If the change was detected in the subinterval $[s^{\ast},e^{\ast}]$ then DAIS will be reapplied to $[1,s^{\ast}]$ and $[e^{\ast},T]$. When checking any interval $[s,e]$, a contrast function is applied to the data, testing for possible change-points at location b, where $b\in\{s,s+1,\ldots,e-1\}$. The contrast function is denoted by $C^{b}_{s,e}(\boldsymbol{X})$ and the maximum value attained in the interval is compared to a threshold $\zeta_{T}$. If $C^{b}_{s,e}(\boldsymbol{X})$ exceeds the threshold, we take $b$ to be a change-point. If no change-point is detected, then the algorithm terminates. The choice of the contrast function for the different cases of signals $f_{t}$ in (1), is explained in Sections 3.1 and 3.2, while the choice of the values for $\zeta_{T}$ and $\lambda_{T}$ is explained in Section 4.2.
@@
There are two main advantages to the algorithm due to its data-adaptive nature. Firstly, since DAIS starts checking for potential change-points in an area where there is reason to believe that includes the location of the change-point with the change of the largest magnitude, the change-point is detected fast. Secondly, when the change-point is at the location of the largest difference, or close to it, due to the way the expansions occur, the change-point lies in the middle of the intervals being checked when an even number of expansions have been performed. This provides advantages in detection power due to the change-point being detected in balanced intervals. Thus, the data-adaptive nature of the algorithm enhances DAIS’ speed and accuracy in estimating the locations of the change-points.
@@ 2.3 Location of the largest difference and the isolation aspect
As already explained in detail in Section 2, the algorithm calculates the location of the largest difference $d_{s,e}$, as defined in (2), in the interval $[s,e]$ under consideration at each step and checks for change-points around it. In this subsection, we explain the role of the location of the largest difference in the detection of a change-point in an interval where isolation is guaranteed.
@@
As will be explained in the proof of Theorem 1 in Appendix B, the detection of the change-point $r_{j}$ will certainly occur if both end-points are at least at a distance $\delta_{T}/2n$ from the true location of the change-point, but it can happen at any smaller interval where the value of the contrast function exceeds the pre-specified threshold $\zeta_{T}$. The positive number $n$ satisfies $3/2\leq n\leq\delta_{T}/2\lambda_{T}$ where, for $r_{0}=1,r_{N+1}=T$, $\delta_{T}=\min_{j=1,\ldots,N+1}\{r_{j}-r_{j-1}\}$ is the smallest distance between two consecutive change-points and $\lambda_{T}$ is the expansion parameter. In the worst-case scenario, meaning that the interval in which detection occurs is the biggest possible and therefore the most expansions have been performed, it holds that, firstly, both end-points are at least a distance $\delta_{T}/2n$ from the change-point and, secondly, at least one of the end-points will be in one of the intervals:
$I^{L}_{j}=\left(r_{j}-\frac{\delta_{T}}{n},r_{j}-\frac{\delta_{T}}{2n}\right],\quad I^{R}_{j}=\left[r_{j}+\frac{\delta_{T}}{2n},r_{j}+\frac{\delta_{T}}{n}\right)$ (5)
@@
for $3/2\leq n\leq\delta_{T}/2\lambda_{T}$. Setting $n=\frac{\delta_{T}}{2\lambda_{T}}$ ensures that the detection happens as quickly as possible. Note that in [2] and [18] $n=3/2$ while in [5] $n=3$ when defining similar intervals for the location of the end-points of the interval where detection occurs.
@@
Since the length of the intervals in (5) is $\delta_{T}/2n$, the condition $2n\lambda_{T}\leq\delta_{T}$ is required, so that there is always at least one end-point that lies within $I^{L}_{j}$ and $I^{R}_{j}$. Since $\lambda_{T}\geq 1$, we implicitly require $\delta_{T}\geq 2n$. Define $\delta_{j}=r_{j+1}-r_{j}$ and $\delta_{s,e}^{j}=\min\{r_{j+1}-d_{s,e},d_{s,e}-r_{j}\}$ for $j=0,1,\ldots,N$. Without loss of generality, suppose that $d_{s,e}\in[r_{J},r_{J+1}]$ for some $J\in\{0,1,\ldots,N\}$. Isolation is guaranteed whenever the following holds
$0\leq\delta_{s,e}^{J}\leq\frac{\delta_{J}}{2}-\frac{3\delta_{T}}{4n}.$ (6)
@@
Note that the lower bound of $n$ ensures that the right hand side of the above equation is non-negative. The bound in (6) both guarantees detection in an interval where the change-point is isolated but also allows for the detection of the neighbouring change-point. This is because the end-point of the interval the algorithm will be applied to after detection of the first change-point (either $r_{J}$ or $r_{J+1}$), will have a distance of at least $\delta_{T}/2n$ from all the undetected change-points; see (B) and (D) in the Appendix for a proof that $\delta_{T}/2n$ is the minimum distance of the end-points from the change-point that guarantees detection.
It should be mentioned that $\delta_{j}=\mathcal{O}(\delta_{T})$ for any $j=0,1,\ldots,N$. Also, since $2n\lambda_{T}\leq\delta_{T}$, $n\lambda_{T}=\mathcal{O}(\delta_{T})$ and the upper bound of $\delta_{s,e}^{J}$ is $\mathcal{O}(\delta_{T})$. Now, as is given by assumption (A2) of Section 3.1, it must hold that $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$. This means that when $T\rightarrow\infty$ then either $\delta_{T}$ or $\underline{f}_{T}$ increases or both do.
  • If $\delta_{T}$ increases then the upper bound of $\delta_{s,e}^{J}$ increases with the same rate.
  • If $\underline{f}_{T}$ increases, the signal-to-noise ratio becomes larger and so the probability that the largest difference is exactly at the location of the change-point increases.
In either case, the probability that $d_{s,e}$ lies within $\delta_{s,e}^{J}$ of any change-point increases as $T$ becomes large. For the case of piecewise-linear signals, we can reach the same conclusion using assumption (A3) in Section 3.2.
@@
$r_{J}-\frac{\delta_{J-1}}{2}+\frac{3\delta_{T}}{4n}\leq d_{s,e}\leq r_{J}+\frac{\delta_{J}}{2}-\frac{3\delta_{T}}{4n}$ (7)
for the change-point $r_{J}$ that is closest to $d_{s,e}$. Note that the length of the interval where $d_{s,e}$ cannot lie under this assumption is $2(\lceil 3\delta_{T}/4n\rceil-1)$ which can be as small as $23\lambda_{T}-$ 2. This means that for appropriately chosen values of $n$ and $\lambda_{T}$, any location is an acceptable location for $d_{s,e}$ besides the midpoint between consecutive change-points.

### v2

@@ 2.2 The DAIS algorithm
In the following paragraphs, the method is presented in more generality. For the rest of the paper, we use $s,e$ to denote the start- and the end-point, respectively, of the interval under consideration, while $T\in\mathbb{N}$ denotes the total length of the given signal and $1\leq s<e\leq T$. We denote by $N\in\mathbb{N}_{0}$ and by $r_{i}$ for $i=1,2,\dots,N$ the number and the locations of the true change-points, respectively, with the latter ones sorted in increasing order, while $\hat{N}$ and $\hat{r}_{i}$ for $i=1,2,\dots,\hat{N}$ are their estimated values. The true, unknown number, $N$, of change-points is allowed to grow with the sample size. We denote by $\lambda_{T}\in\mathbb{N}$ the expansion parameter and $\zeta_{T}\in\mathbb{R^{+}}$ the threshold, which is used to decide whether a time-point is a change-point; more details on the choice of their values are given in Section 4.2. In addition, $d_{s,e}$ is the location of the largest difference, in absolute value, in the interval $[s,e]$ and is defined as
@@
We denote by $C^{b}_{s,e}(\boldsymbol{X})$ the contrast function applied to the location $b$ using the values $\{X_{s},X_{s+1},\ldots,X_{e}\}$ for $b\in[s,e)$. The functions used for the cases of piecewise-constant and continuous, piecewise-linear signals are explained in Sections 3.1 and 3.2, respectively. For $K^{l}=\lceil\frac{d_{s,e}-s+1}{\lambda_{T}}\rceil$, $K^{r}=\lceil\frac{e-d_{s,e}+1}{\lambda_{T}}\rceil$, $K^{\max}=\max\{K^{l},K^{r}\}$, $K^{\min}=\min\{K^{l},K^{r}\}$, and $K=K^{\max}+K^{\min}$, then the start- and end-points of the intervals $[c_{m}^{l},c_{k}^{r}]$ in which the algorithm tests for the existence of a change-point, are given by
@@
These are collected in the ordered sets, $L_{d_{s,e},s,e}$ and $R_{d_{s,e},s,e}$, defined in (2.2), which consist of the left and right end-points, respectively, of the intervals used in DAIS. Since the expansions occur in turn once from the right and once from the left, the points appear twice in each set, up until the moment when one end-point is equal to either $s$ or $e$. The left- and right-expanding intervals checked are collected in the ordered set $I_{j}$ in (4). The ordered sets of the end-points for the expansions are:
$L_{d_{s,e},s,e}=\bigl{\{}c^{l}_{0},c^{l}_{1},c^{l}_{1},c^{l}_{2},\dots,c^{l}_{K^{\min}-1},c^{l}_{K^{\min}},c^{l}_{K^{\min}+1},c^{l}_{K^{\min}+2},\ldots,c^{l}_{K^{\max}}\bigr{\}},$
$R_{d_{s,e},s,e}=\bigl{\{}c^{r}_{1},c^{r}_{1},c^{r}_{2},c^{r}_{2},\dots,c^{r}_{K^{\min}},c^{r}_{K^{\min}},c^{r}_{K^{\min}+1},c^{r}_{K^{\min}+2},\ldots,c^{r}_{K^{\max}}\bigr{\}},$ (3)
@@
$I_{j}=\bigl{[}s_{j},e_{j}\bigr{]}=\Bigl{[}L_{d_{s,e},s,e}[j],R_{d_{s,e},s,e}[j]\Bigr{]},\textrm{ for }j\in\{1,2,\ldots,K\},$ (4)
where A[i] denotes the $i^{th}$ element of the ordered set A. It is easy to show that $\lvert L_{d_{s,e},s,e}\rvert=\lvert R_{d_{s,e},s,e}\rvert=K$, where $\lvert A\rvert$ denotes the cardinality of the set A. Note that the largest number of expansions required in order to check the whole length of the interval $[s,e]$ around $d_{s,e}$ is $\lceil\frac{e-s+1}{\lambda_{T}}\rceil+1$. The intervals $I_{j}$ defined in (4) are not deterministic. The end-points depend on the location of the largest difference, which depends on the structure of the unknown signal and the noise of the data sequence under consideration.
@@
     $b_{i}=\rm{argmax}_{t\in[s_{i},e_{i})}C_{s_{i},e_{i}}^{t}\left(\boldsymbol{X}\right)$
     if $C_{s_{i},e_{i}}^{b_{i}}\left(\boldsymbol{X}\right)>\zeta_{T}$ then
@@
Algorithm 1 provides a pseudocode that briefly explains DAIS’ workflow. The first step of DAIS is to identify, within the observed data, $X_{t}$, $t=1,2,\ldots,T$, the location of the largest difference $d_{1,T}$, as defined in (2). This is the step that makes the algorithm data-adaptive, differentiating it from the competitors, as the algorithm will check different intervals $I_{j}$ for the same underlying signals (in terms of the type of the change, the length of the sequence and the change-point locations) if they have different observed noise, depending on the value of $d_{1,T}$. Applying the appropriate transformation (as in (2) based on the signal structure) to the unobserved true signal $f_{t}$ in (1), will reveal a constant signal at 0, with non-zero values at the locations where the change-points occur, so we expect that the largest difference for the observed data $X_{t}$ will be at a point near the change-point with the largest magnitude of change. A proper explanation of what we consider to be ‘near’ is discussed in Section 2.3 in detail. After calculating $d_{1,T}$, expanding intervals are used to check for possible change-points. The algorithm expands the intervals once from the right and once from the left of the point where the maximum absolute difference occurs at every step by an expansion parameter of magnitude $\lambda_{T}$. By doing this, we identify the location of the change-point that is closest to the detected largest difference. This expansion is performed for at most $\lceil\frac{T}{\lambda_{T}}\rceil+1$ steps in total. Upon the detection of a change-point, the algorithm restarts after discarding the data that have already been checked. If, for example, the change-point was detected in the subinterval $[s^{\ast},e^{\ast}]$, then DAIS will be reapplied to $[1,s^{\ast}]$ and $[e^{\ast},T]$. When checking any interval $[s,e]$ the maximum value attained by the contrast function at a point $b$ within the interval, is compared to the chosen threshold $\zeta_{T}$. If this value exceeds $\zeta_{T}$, we take $b$ to be a change-point, for $b\in\{s,s+1,\ldots,e-1\}$ in the case of piecewise-constant signals, and $b\in\{s,s+1,\ldots,e-2\}$ in the case of piecewise-linear signals. If no change-point is detected, then the algorithm terminates.
As mentioned in the introduction, the algorithm can be applied to more general signals, for example piecewise-polynomial signals of order $p$. In this case, the location of the largest difference needs to be generalised. In the more general case, it can be defined as the value of $t$ for which the $(p+1)-$times differentiated signal is maximised.
@@
There are two main advantages to the algorithm due to its data-adaptive nature. Firstly, since at each step DAIS starts checking for potential change-points around the largest difference, which as already explained we have reason to believe that it is around the most prominent change-point, then this change-point is detected fast. If the starting point is not near the change-point, detection still occurs with a slower speed. Secondly, when the change-point is at the location of the largest difference, or close to it, due to the way the expansions occur, the change-point lies in the middle of the intervals being checked when an even number of expansions have been performed. This provides advantages in detection power due to the change-point being detected in some cases in balanced intervals. Such near balance in the distances of the change-points from the left and right end-points of the intervals under consideration does not necessarily appear in many state-of-the-art competitors, such as Binary Segmentation, ID, NOT, WBS, WBS2 and SeedBS. Thus, the data-adaptive nature of the algorithm enhances DAIS’ speed and accuracy in estimating the locations of the change-points.
@@ 2.3 Location of the largest difference and the isolation aspect
As already explained in detail in Section 2.2, for a given data sequence $X_{t}$, the algorithm calculates the location of the largest difference $d_{s,e}$, as defined in (2), in the interval $[s,e]$ under consideration at each step and checks for change-points around it. For different data sequences, DAIS will start searching for change-points in different neighbourhoods of the data, based on where the most prominent change-point (if any) lies. This is exactly the novel data-adaptive nature of the proposed algorithm, compared to the existing algorithms in the literature that exhibit a fixed workflow, not adaptive to the data structure. In this subsection, we now explain the role of the location of the largest difference in the detection of a change-point in an interval where isolation is guaranteed.
@@
As will be explained in the proof of Theorem 1 in Appendix C, the detection of a change-point $r_{j}$, for $j\in\{1,\ldots,N\},N\in\mathbb{N}$ will certainly occur if both end-points are at least at a distance $\delta_{T}/2n$ from the true location of the change-point, but of course it can happen at any smaller interval where the value of the contrast function exceeds the pre-specified threshold $\zeta_{T}$. The positive number $n$ satisfies $3/2\leq n\leq\delta_{T}/2\lambda_{T}$ where, for $r_{0}=0$ and $r_{N+1}=T$, $\delta_{T}=\min_{j=1,\ldots,N+1}\{r_{j}-r_{j-1}\}$ is the smallest distance between two consecutive change-points and $\lambda_{T}$, as explained previously, is the expansion parameter. In the worst-case scenario, meaning that the interval in which detection occurs is the biggest possible and therefore the most expansions have been performed, it holds that, firstly, both end-points are at least a distance $\delta_{T}/2n$ from the change-point and, secondly, at least one of the end-points will be in one of the intervals:
$I^{L}_{j}=\left(r_{j}-\frac{\delta_{T}}{n},r_{j}-\frac{\delta_{T}}{2n}\right],\quad I^{R}_{j}=\left[r_{j}+\frac{\delta_{T}}{2n},r_{j}+\frac{\delta_{T}}{n}\right).$ (5)
@@
Setting $n=\frac{\delta_{T}}{2\lambda_{T}}$ ensures that the detection happens as quickly as possible. Note that in [2] and [23] the choice $n=3/2$ was made while in [6] the value of $n=3$ was used when defining similar intervals for the location of the end-points of the interval where detection occurs.
@@
Since the length of the intervals in (5) is $\delta_{T}/2n$, the condition $2n\lambda_{T}\leq\delta_{T}$ is required, so that there is always at least one end-point that lies within $I^{L}_{j}$ and $I^{R}_{j}$. Since $\lambda_{T}\geq 1$, we implicitly require $\delta_{T}\geq 2n$. For $d_{s,e}\in[1,r_{1}]$ and $d_{s,e}\in[r_{N},T]$, any value of $d_{s,e}$ guarantees isolation, as the left and right expanding intervals will not include a second change-point before detection occurs, besides $r_{1}$ and $r_{N}$, respectively. Now considering the more complicated case, where the location of the largest difference lies between two change-points, define $\tilde{\delta}_{j}=r_{j+1}-r_{j}$ and $\delta_{s,e}^{j}=\min\{r_{j+1}-d_{s,e},d_{s,e}-r_{j}\}$ for $j=0,1,\ldots,N$. Without loss of generality, suppose that $d_{s,e}\in\left(r_{J},r_{J+1}\right)$ for some $J\in\{1,\ldots,N-1\}$. Isolation is guaranteed whenever the following holds
$0\leq\delta_{s,e}^{J}\leq\frac{\tilde{\delta}_{J}}{2}-\frac{3\delta_{T}}{4n}.$ (6)
@@
Note that the lower bound of $n$ ensures that the right hand side of the second inequality in (6) is non-negative. The result in (6) both guarantees detection in an interval where the change-point is isolated but also allows for the detection of the neighbouring change-point. This is because the end-point of the interval the algorithm will be applied to after detection of the first change-point (either $r_{J}$ or $r_{J+1}$), will have a distance of at least $\delta_{T}/2n$ from all the undetected change-points; see (C) and (D) in the Appendix for a proof that $\delta_{T}/2n$ is the minimum distance of the end-points from the change-point that guarantees detection.
It should be mentioned that $\tilde{\delta}_{j}=\mathcal{O}(\delta_{T})$ for any $j=0,1,\ldots,N$. Also, since $2n\lambda_{T}\leq\delta_{T}$, $n\lambda_{T}=\mathcal{O}(\delta_{T})$ and, from (6), the upper bound of $\delta_{s,e}^{J}$ is $\mathcal{O}(\delta_{T})$. Now, as is given by assumption (A2) of Section 3.1, it must hold that $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$, where $\underline{f}_{T}=\min_{j}|f_{r_{j}+1}-f_{r_{j}}|$ is defined as the minimum magnitude of jumps. This assumption implies that when $T\rightarrow\infty$ then $\sqrt{\delta_{T}}\underline{f}_{T}$ increases. If $\delta_{T}$ increases then the upper bound of $\delta_{s,e}^{J}$ increases with the same rate. If $\underline{f}_{T}$ increases, the signal-to-noise ratio becomes larger and so the probability that the largest difference is exactly at the location of the change-point increases. More details on this, in the case that $\epsilon_{t}$ is Gaussian, are provided in Section 3.3. In either case, the probability that $d_{s,e}$ lies within $\delta_{s,e}^{J}$ of any change-point increases as $T$ becomes large. We can reach the same conclusions using Assumptions (A3) and (A4) for the case of univariate piecewise-linear signals and multivariate signals, respectively.
@@
$d_{s,e}\in\left[1,r_{1}+\frac{\tilde{\delta}_{1}}{2}-\frac{3\delta_{T}}{4n}\right]\cup\left[r_{N}-\frac{\tilde{\delta}_{N-1}}{2}+\frac{3\delta_{T}}{4n},T\right]\cup\left\{\bigcup_{j=2}^{N-1}B_{n;j}\right\}$ (7)
where
$B_{n;j}=\left[r_{j}-\frac{\tilde{\delta}_{j-1}}{2}+\frac{3\delta_{T}}{4n},r_{j}+\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}\right].$ (8)
Note that the length of the interval where $d_{s,e}$ cannot lie under this assumption is $2(\lceil 3\delta_{T}/4n\rceil-1)$ around the midpoint of two consecutive change-points. This can be as small as $23\lambda_{T}-$ 2, which means that for appropriately chosen values of $n$ and $\lambda_{T}$, any location is an acceptable location for $d_{s,e}$ besides the midpoint between consecutive change-points.